OpenStudy (anonymous):

what is the molar mass of 4 moles of Cu(CN)2

OpenStudy (anonymous):

ANYBODY THERE

OpenStudy (anonymous):

find the atomic mass of each element then add them together and multiply them by 4

OpenStudy (anonymous):

Cu(63.5)+Cx2(12.01x2=24.02)+Nx2(14.01x2=28.02) = 115.54 molar mass of one mole. Then multiply by 4, the number of moles. 115.54x4=462.16